OpenStudy (anonymous):

what is the equation of the line passing through (–6, 1) and perpendicular to the line y = –3x + 1 in slope-intercept form

OpenStudy (anonymous):

Get the slope of the line y = -3x + 1, which is in the slope intercept form y = mx + c, Where m is the slope. Invert that slope and change the sign to get the slope of any line perpendicular to it. Then use the equation of the line formula \[y - y _{1} = m(x - x _{1})\] Where m is the new slope you found, And \[x _{1} and y _{1}\] are points on the line for which the equation you need to find.

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):

yes, thank you c:

OpenStudy (anonymous):

glad to help